Simplify the expression.
[4 - 7i] + [2 + 3i]

Answer by stanbon(75887)

Simplify the expression.
[4 - 7i] + [2 + 3i]
-----
= 8 + 12i - 14i -21i^2
----------
= 8 + 21 - 2i
-----
= 29 - 2i
-----------
Cheers,
Stan H.
===========